OpenStudy (anonymous):

Use scientific notation to multiply the following two numbers. Express the answer using scientific notation; retain all nonzero decimal places. (1.6 · 10^4) · (9.9 · 10^6)

OpenStudy (paxpolaris):

What did u get?

OpenStudy (anonymous):

nothing i don't understand it

OpenStudy (paxpolaris):

\[(1.6 · 10^4) \times (9.9 · 10^6) =(1.6 \cdot 9.9)\times(10^4\cdot 10^6)\]

OpenStudy (paxpolaris):

\[\large=1.6\times 9.9 \times (10^{(6+4)})= (1.6\times9.9)\cdot 10^{10}\]

OpenStudy (paxpolaris):

\[\large = 15.84 \cdot 10^{10}=\huge1.584 \cdot 10^{11}\]

OpenStudy (anonymous):

is that the whole thing

OpenStudy (paxpolaris):

yes... do you understand scientific notation?